# Establishing Linear Surrogate Regret Bounds for Convex Smooth Losses via Convolutional Fenchel–Young Losses

### v2

@@ A.4 Proof of Lemma 14
$\underline{R}_{L_{\Omega_{T}}}(\bm{\eta})$ $=\inf_{\bm{\theta}\in\mathbb{R}^{\varrho}}R_{L_{\Omega_{T}}}(\bm{\theta},\bm{\eta})$
$=\inf_{\bm{\theta\in\mathbb{R}^{\varrho}}}\Bigl{[}\Omega_{T}^{*}(\bm{\theta})-\langle\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\rangle\Bigr{]}+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]$
$=-\sup_{\bm{\theta\in\mathbb{R}^{\varrho}}}\Bigl{[}\langle\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\rangle-\Omega_{T}^{*}(\bm{\theta})\Bigr{]}+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]$
@@
=(A)${=}\underbrace{\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Bigl{\langle}\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{\rangle}+\Omega\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}}_{\geq 0\quad\text{due to the Fenchel--Young inequality}}$
@@
(B)${\geq}\sum_{t=1}^{N}\pi_{t}\Bigl{\langle}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)],\bm{\ell}^{\bm{\rho}}(t)\Bigr{\rangle}-\min_{t\in\widehat{\mathcal{Y}}}\Bigl{\langle}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)],\bm{\ell}^{\bm{\rho}}(t)\Bigr{\rangle}$
=(C)${=}\sum_{t=1}^{N}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta}),$
@@
A.5 Proof of Theorem 17
Proof.  First of all, we prove that $\mathrm{relint}(\mathrm{conv}(\bm{\rho}(\mathcal{Y})))$ is in the image of $\nabla\Omega_{T}^{*}$ by contradiction. According to Rockafellar [57, Theorem 23.4], we have that $\partial\Omega_{T}(\bm{p})$ is non-empty for all $\bm{p}\in\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$. Now suppose there exists $\bm{p}\in\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ such that $\bm{p}\not=\nabla\Omega_{T}^{*}(\bm{\theta})$ for any $\bm{\theta}\in\mathbb{R}^{\varrho}$. Since $\partial\Omega_{T}(\bm{p})$ is non-empty, there exists $\bm{\theta}^{\prime}\in\mathbb{R}^{\varrho}$ such that $\bm{\theta}^{\prime}=\nabla\Omega_{T}(\bm{p})$. By Rockafellar and Wets [58, Proposition 11.3] (applied on proper l.s.c. $\Omega_{T}$), we have $\bm{p}=\nabla\Omega_{T}^{*}(\bm{\theta^{\prime}})$, which contradicts the assumption $\bm{p}\not=\nabla\Omega_{T}^{*}(\bm{\theta}^{\prime})$. Thus we have verified that $\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ is in the image of $\nabla\Omega_{T}^{*}$. This additionally implies that $\mathrm{relint}(\mathrm{conv}(\bm{\rho}(\mathcal{Y})))\subseteq\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ is also in the image of $\nabla\Omega_{T}^{*}$ because of Condition 1.
@@ A.5 Proof of Theorem 17
is differentiable and convex in $\bm{\theta}$. If $\bm{\theta}^{*}\in\mathbb{R}^{\varrho}$ is the minimizer of $R_{L_{\Omega_{T}}}(\cdot,\bm{\eta})$, its optimality condition indicates
@@
Appendix B Detailed Proofs in Section 4
B.1 Proof of Corollary 18
@@
where $\alpha_{1},\dots,\alpha_{K}\geq 0$ and $\beta\geq 0$ are the Lagrangian multipliers. Since $V_{\bm{\theta}}(\bm{\pi})$ is convex and differentiable, and the feasible region $\Delta^{K}$ is convex, the KKT conditions are necessary and sufficient for the optimality of $\bm{\pi}^{*}\in\Pi(\bm{\theta})$, which requires that there exists $(\bm{\alpha}^{*},\beta^{*})$ satisfying
$-e^{\theta_{y}+1-\pi_{y}^{*}}-\alpha_{i}^{*}+\beta^{*}=0,$ for any $y=1,\dots,K$, (20)
@@
Denote by $f(\tau)\coloneqq\sum_{y=1}^{K}\max\{\theta_{y}-\tau,0\}$. Then $f$ is continuous on $\mathbb{R}$, and moreover, it is strictly decreasing on $(-\infty,\theta_{[1]}]$ and equals 00 for $\tau\geq\theta_{[1]}$ because of the following expression:
$f(\tau)=\begin{cases}\sum_{y=1}^{K}(\theta_{y}-\tau)&\text{if $\tau\leq\theta_{[K]}$,}\\ \sum_{i=1}^{k-1}(\theta_{[i]}-\tau)&\text{if $\tau\in[\theta_{[k]},\theta_{[k-1]}]$ for $k=2,\dots,K$,}\\ 0&\text{if $\tau>\theta_{[1]}$}.\end{cases}$
Then $n$ defined in line 2 of Algorithm 1 exists since $f(\theta_{[1]})=0<1$. We also have that $f(\theta_{[n]})=\sum_{i=1}^{n}(\theta_{[i]}-\theta_{[n]})<1$. When $n<K$, we have that $f(\theta_{[n+1]})=\sum_{i=1}^{n+1}(\theta_{[i]}-\theta_{[n+1]})=\sum_{i=1}^{n}(\theta_{[i]}-\theta_{[n+1]})\geq 1$ by definition. Then we see
$\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}\in[\theta_{[n+1]},\theta_{[n]}]$
@@
$f\left(\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}\right)$ $=\sum_{i=1}^{n}\left(\theta_{[i]}-\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}\right)$
$=\sum_{i=1}^{n}\theta_{[i]}-\sum_{i=1}^{n}\theta_{[i]}+1$
@@
When $n=K$, we have that $\sum_{i=1}^{K}\theta_{[i]}-K\theta_{[k]}<1$, that is,
$\frac{\sum_{i=1}^{K}\theta_{[i]}-1}{K}<\theta_{[K]}.$
@@
$f\left(\frac{\sum_{i=1}^{K}\theta_{[i]}-1}{K}\right)$ $=\sum_{i=1}^{K}\left(\theta_{[i]}-\frac{\sum_{i=1}^{K}\theta_{[i]}-1}{K}\right)$
$=\sum_{i=1}^{K}\theta_{[i]}-\sum_{i=1}^{K}\theta_{[i]}+1$
@@
$\tau^{*}=\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}<\theta_{[1]},$
which is what line 3 of Algorithm 1 returns. Since $f(\tau)$ is strictly decreasing on $(-\infty,\theta_{[1]}]$, $\tau^{*}$ uniquely satisfies $f(\tau^{*})=1$.
Since $\sum_{i=1}^{k}\theta_{[i]}$ can be calculated cumulatively in $\mathcal{O}(K)$ time and the quick sort runs in $\mathcal{O}(K\log K)$ time, Algorithm 1 runs in $\mathcal{O}(K\log K)$ time in total. Furthermore, since $\tau^{*}$ is unique and $\pi_{y}^{*}=\max\{\theta_{y}-\tau^{*},0\}$, we ensure that $\Pi(\bm{\theta})$ is a singleton, which concludes the proof.
B.2 Proof of Proposition 19
@@
$\mathop{\mathrm{missing}}{argmax}_{t\in\mathcal{Y}}\theta_{t}=\mathop{\mathrm{missing}}{argmax}_{t\in\mathcal{Y}}{{\pi}}_{\log}(\bm{\theta})_{t}.$
According to Corollary 18 and Algorithm 1, there exists $\tau\in\mathbb{R}$ such that $\pi_{\log}(\theta)_{i}=\max\{\theta_{i}-\tau,0\}$. We also have that $\max_{t}\theta_{t}>\tau$ by contradiction: if $\max_{t}\theta_{t}\leq\tau$, $\bm{\pi}_{\log}(\bm{\theta})=\bm{0}$ holds, which contradicts $\bm{\pi}_{\log}(\bm{\theta})\in\Delta^{K}$.

### v3

@@ B.5 Proof of Lemma 14
$\underline{R}_{L_{\Omega_{T}}}(\bm{\eta})$ $=\inf_{\bm{\theta}\in\mathbb{R}^{d}}R_{L_{\Omega_{T}}}(\bm{\theta},\bm{\eta})$
$=\inf_{\bm{\theta\in\mathbb{R}^{d}}}\Bigl[\Omega_{T}^{*}(\bm{\theta})-\langle\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\rangle\Bigr]+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]$
$=-\sup_{\bm{\theta\in\mathbb{R}^{d}}}\Bigl[\langle\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\rangle-\Omega_{T}^{*}(\bm{\theta})\Bigr]+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]$
@@
$\overset{\text{(A)}}{=}\underbrace{\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Bigl\langle\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr\rangle+\Omega\Bigl(\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr)}_{\geq 0\quad\text{due to the Fenchel--Young inequality}}$
@@
$\overset{\text{(B)}}{\geq}\sum_{t=1}^{N}\pi_{t}\Bigl\langle\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)],\bm{\ell}^{\bm{\rho}}(t)\Bigr\rangle-\min_{t\in\widehat{\mathcal{Y}}}\Bigl\langle\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)],\bm{\ell}^{\bm{\rho}}(t)\Bigr\rangle$
$\overset{\text{(C)}}{=}\sum_{t=1}^{N}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta}),$
@@
B.6 Proof of Corollary 16
[…]
Remark 21
Interestingly, rama_cvxdim also used the affine dimension of the loss matrix to study the dimension of convex surrogates, suggesting that this concept plays a important role in loss function design and deserves more attention.
B.7 Proof of Theorem 17
Proof. First of all, we prove that $\mathrm{relint}(\mathrm{conv}(\bm{\rho}(\mathcal{Y})))$ is in the image of $\nabla\Omega_{T}^{*}$ by contradiction. According to convex, we have that $\partial\Omega_{T}(\bm{p})$ is non-empty for all $\bm{p}\in\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$. Now suppose there exists $\bm{p}\in\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ such that $\bm{p}\not=\nabla\Omega_{T}^{*}(\bm{\theta})$ for any $\bm{\theta}\in\mathbb{R}^{d}$. Since $\partial\Omega_{T}(\bm{p})$ is non-empty, there exists $\bm{\theta}^{\prime}\in\mathbb{R}^{d}$ such that $\bm{\theta}^{\prime}=\nabla\Omega_{T}(\bm{p})$. By variational (applied on proper l.s.c. $\Omega_{T}$), we have $\bm{p}=\nabla\Omega_{T}^{*}(\bm{\theta^{\prime}})$, which contradicts the assumption $\bm{p}\not=\nabla\Omega_{T}^{*}(\bm{\theta}^{\prime})$. Thus we have verified that $\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ is in the image of $\nabla\Omega_{T}^{*}$. This additionally implies that $\mathrm{relint}(\mathrm{conv}(\bm{\rho}(\mathcal{Y})))\subseteq\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ is also in the image of $\nabla\Omega_{T}^{*}$ because of Condition 1.
@@ B.7 Proof of Theorem 17
is differentiable and convex in $\bm{\theta}$. If $\bm{\theta}^{*}\in\mathbb{R}^{d}$ is the minimizer of $R_{L_{\Omega_{T}}}(\cdot,\bm{\eta})$, its optimality condition indicates
@@
Appendix C Deferred Proofs in Section 4
C.1 Proof of Lemma 18
@@
where $\alpha_{1},\dots,\alpha_{K}\geq 0$ and $\beta$ are the Lagrangian multipliers. Since $V_{\bm{\theta}}(\bm{\pi})$ is convex and differentiable, and the feasible region $\Delta^{K}$ is convex, the KKT conditions are necessary and sufficient for the optimality of $\bm{\pi}^{*}\in\Pi(\bm{\theta})$, which requires that there exists $(\bm{\alpha}^{*},\beta^{*})$ satisfying
$-e^{\theta_{y}+1-\pi_{y}^{*}}-\alpha_{y}^{*}+\beta^{*}=0,$ for any $y=1,\dots,K$, (20)
@@
Denote by $f(\tau)\coloneqq\sum_{y=1}^{K}\max\{\theta_{y}-\tau,0\}$. Then $f$ is continuous on $\mathbb{R}$, and moreover, it is strictly decreasing on $(-\infty,\theta_{(1)}]$ and equals 0 for $\tau\geq\theta_{(1)}$ because of the following expression:
$f(\tau)=\begin{cases}\sum_{y=1}^{K}(\theta_{y}-\tau)&\text{if $\tau\leq\theta_{(K)}$,}\\ \sum_{i=1}^{k-1}(\theta_{(i)}-\tau)&\text{if $\tau\in[\theta_{(k)},\theta_{(k-1)}]$ for $k=2,\dots,K$,}\\ 0&\text{if $\tau>\theta_{(1)}$}.\end{cases}$
Then $n$ defined in line 2 of Algorithm 1 exists since $f(\theta_{(1)})=0<1$. We also have that $f(\theta_{(n)})=\sum_{i=1}^{n}(\theta_{(i)}-\theta_{(n)})<1$. When $n<K$, we have that $f(\theta_{(n+1)})=\sum_{i=1}^{n+1}(\theta_{(i)}-\theta_{(n+1)})=\sum_{i=1}^{n}(\theta_{(i)}-\theta_{(n+1)})\geq 1$ by definition. Then we see
$\frac{\sum_{i=1}^{n}\theta_{(i)}-1}{n}\in[\theta_{(n+1)},\theta_{(n)}]$
@@
$f\left(\frac{\sum_{i=1}^{n}\theta_{(i)}-1}{n}\right)$ $=\sum_{i=1}^{n}\left(\theta_{(i)}-\frac{\sum_{i=1}^{n}\theta_{(i)}-1}{n}\right)$
$=\sum_{i=1}^{n}\theta_{(i)}-\sum_{i=1}^{n}\theta_{(i)}+1$
@@
When $n=K$, we have that $\sum_{i=1}^{K}\theta_{(i)}-K\theta_{(k)}<1$, that is,
$\frac{\sum_{i=1}^{K}\theta_{(i)}-1}{K}<\theta_{[K]}.$
@@
$f\left(\frac{\sum_{i=1}^{K}\theta_{(i)}-1}{K}\right)$ $=\sum_{i=1}^{K}\left(\theta_{(i)}-\frac{\sum_{i=1}^{K}\theta_{(i)}-1}{K}\right)$
$=\sum_{i=1}^{K}\theta_{(i)}-\sum_{i=1}^{K}\theta_{(i)}+1$
@@
$\tau^{*}=\frac{\sum_{i=1}^{n}\theta_{(i)}-1}{n}<\theta_{(1)},$
which is what line 3 of Algorithm 1 returns. Since $f(\tau)$ is strictly decreasing on $(-\infty,\theta_{(1)}]$, $\tau^{*}$ uniquely satisfies $f(\tau^{*})=1$.
Since $\sum_{i=1}^{k}\theta_{(i)}$ can be calculated cumulatively in $\mathcal{O}(K)$ time and the quick sort runs in $\mathcal{O}(K\log K)$ time, Algorithm 1 runs in $\mathcal{O}(K\log K)$ time in total. Furthermore, since $\tau^{*}$ is unique and $\pi_{y}^{*}=\max\{\theta_{y}-\tau^{*},0\}$, we ensure that $\Pi(\bm{\theta})$ is a singleton, which concludes the proof.
C.2 Proof of Proposition 19
@@
$\mathop{\mathrm{argmax}}_{t\in\mathcal{Y}}\theta_{t}=\mathop{\mathrm{argmax}}_{t\in\mathcal{Y}}{{\pi}}_{\log}(\bm{\theta})_{t}.$
According to Lemma 18 and Algorithm 1, there exists $\tau\in\mathbb{R}$ such that $\pi_{\log}(\theta)_{i}=\max\{\theta_{i}-\tau,0\}$. We also have that $\max_{t}\theta_{t}>\tau$ by contradiction: if $\max_{t}\theta_{t}\leq\tau$, $\bm{\pi}_{\log}(\bm{\theta})=\bm{0}$ holds, which contradicts $\bm{\pi}_{\log}(\bm{\theta})\in\Delta^{K}$.